OpenStudy (gabylovesyou):

why is x^2 - 16 --------- x^2 - 4x a perfect square ?

OpenStudy (luigi0210):

because you can separate it into two equations with the same and opposite number..? I don't know how to explain this ^_^' (x-4)(x+4)

OpenStudy (gabylovesyou):

but why 4 ?

OpenStudy (luigi0210):

it multiplies to 16

OpenStudy (gabylovesyou):

ohhhhhhhh ok

OpenStudy (luigi0210):

Not really a good explanation but it's all I got :P

OpenStudy (campbell_st):

I don't think its a perfect square as it factorises to \[\frac{(x + 4)(x -4)}{x(x - 4)}\]

OpenStudy (campbell_st):

just a thought

OpenStudy (luigi0210):

>.< I did not see that fraction there..

OpenStudy (gabylovesyou):

how do i do 3x^2 + 15

OpenStudy (luigi0210):

factor out a 3

OpenStudy (campbell_st):

a perfect square \[(a + b)^2 = (a + b)(a + b) = a^2 + 2ab + b^2\] I always thought they has 3 terms... your question has the difference of 2 squares in the numerator and a quadratic binomial in the denominator

OpenStudy (gabylovesyou):

u cant its just 3 isnt it ?

OpenStudy (gabylovesyou):

ohh ok @campbell_st

OpenStudy (luigi0210):

3(x^2+5)

OpenStudy (luigi0210):

wait are you solving it?

OpenStudy (gabylovesyou):

howd u get that :/

OpenStudy (gabylovesyou):

u cant do anything with the 3 so it stays like that .. x^2 , one will go with the 3 (3x) and the other in parenthesis 3x(x but howd u get + 5 ? cuz of 3 * 5 ? 3x(x + 5)

OpenStudy (luigi0210):

Yup

OpenStudy (gabylovesyou):

k thanks